Calculate the quantities for ideal gas given below.
(R=0.082 L atm mol-1 K-1)

iv. A closed gas cylinder contains 0.500 mole H2, 0.300 mole O2, 1.200 mole CO2 at 25 °C and pressures 2.00 atm. Calculate the volume of the cylinder.
v. Calculate the partial pressure of O2 inside the cylinder in question (iv).
1
Expert's answer
2020-07-22T07:59:41-0400

Question (iv)

The number of moles of hydrogen, oxygen, and carbondioxide are given.

Temperature = 25oC which is equal to 298K

The total pressure inside the cylinder is equal to summation of pressures given = 2.0 atm

But PV=nRT then rewtiting the equation:

V=(nRT)/P

Also, the number of particles present that hit the walls of the cylinder and hence the pressure can be expressed in terms of the total number of moles of all the gases = 1.2+0.3+0.5 = 2.0 atm

V= (2*0.82*298)/2

Volume of the cylinder = 24.44L

Question (v)

Partial pressure of oxygen = (nRT)/V

P = (2*0.82*298)/24.44

P = 0.3atm